# Anchors Aweigh! Sail for Optimal Unified Multi-Modal Representations

### v1

@@ 4 Experiment
4.1 Experiments with synthetic dataset
Synthetic datasets.
We employ a latent variable model (Bishop & Nasrabadi, 2006) for generating synthetic multimodal datasets. A latent variable model is a statistical model for data ${\bf X}\in\mathbb{R}^{d_{x}}$, under which ${\bf X}$ is generated according to a conditional probability distribution $P_{{\bf X}|{\bf Z}}$, where ${\bf Z}\in\mathbb{R}^{d_{z}}$ is the latent variable. In terms of the representation learning framework, ${\bf Z}$ can be seen as the true representation of ${\bf X}$. We assume that the class label ${\bf Y}\in[K]$ and the latent variable ${\bf Z}$ are jointly distributed according to $P_{{\bf Z},{\bf Y}}$. In our setting, we exploit Gaussian mixture model (GMM) (Bishop & Nasrabadi, 2006) for the latent variable ${\bf Z}$, and we generate $M$ modalities ${\bf X}_{i}=g_{i}({\bf Z})+{\bf N},i\in[M]$ with random noise ${\bf N}$ and some non-linear projections $g_{i}:\mathbb{R}^{d_{z}}\to\mathbb{R}^{d_{x}}$. We choose the projections in a way such that each model can be ranked in ascending order, i.e., ${\bf X}_{1}$ is the worst, and ${\bf X}_{4}$ is the best modality in terms of their inherent correlation with the latent variable. The class label ${\bf Y}$ is set to the component id of GMM (for details, see Appendix C.1).
@@ 4.1 Experiments with synthetic dataset
Figure 2 shows the classification accuracies with a synthetic dataset of $M=4$ modalities. To obtain the results in Figure 2(a), we initialize pretrained backbones for all modalities, apply FABind (${\bf X}_{i}$-B) with anchor ${\bf X}_{i}$ or CentroBind (CB), and evaluate accuracy (acc(${\bf Z}_{i}$)) with embeddings from $i$-th modality. We provide acc(${\bf Z}_{i}$), without any binding, for a reference. Figure 2(a) verifies our analysis of FABind (Section 2.3) and CentroBind (Section 3.2): (1) the comparison between ${\bf X}_{1}$-B and ${\bf X}_{4}$-B shows the importance of choosing anchor modality; (2) the comparison between acc(${\bf Z}_{4}$) and ${\bf X}_{1}$-B: acc(${\bf Z}_{4}$) shows a performance deterioration by FABind , demonstrating that FABind does not capture intra information; (3) CB consistently outperform all FABind , indicating that CB successfully captures elements that FABind overlooks, including intra information and shared information among non-anchored modalities.
Figure 2(b) includes accuracies of FABind and CB with random backbones. Similarly, CB outperforms all baselines. Somewhat surprisingly, CB with random backbones (green curves) also performs better than FABind with pretrained backbones (red curves). This further supports our analysis that CentroBind is robust to the backbone quality as it optimizes intra and shared information, whereas FABind is sensitive to the backbone quality. Overall, these empirical results validate our findings. We provide additional experimental results on synthetic datasets with $M=6,8$ in Appendix C.1. With the larger number of modalities, CB still outperforms the baselines, strengthening our analysis of CB.
@@ 4.2 Experiments with real-world dataset
In this section, we provide experiment results with a real-world dataset. We compare CentroBind and FABind anchored at text modality. We utilize the MUStARD dataset  (Castro et al., 2019) for its rich combination of multimodal data with more than two modalities. It consists of 690 video clips (including audio) and text for sarcasm detection with labels such as sarcasm indicators and speaker names. For the backbones in FABind and CentroBind , we use the pretrained VideoMAE model (Tong et al., 2022) for video data, the pretrained WaveLM model (Chen et al., 2022) for audio data, and the pretrained BERT model (Devlin et al., 2019) for text data. A detailed description of the training setting is provided in Appendix C.2.
@@
We perform evaluations in zero-shot binary and multi-class classification tasks, One-to-One cross-modal retrieval, and Two-to-One cross-modal retrieval. For classification tasks, we use a Multi-Layer Perceptron (MLP) to perform sarcasm detection as a binary classification and speaker classification with $23$ multi-class categories. In particular, MLP is trained on embeddings in a single modality (denoted by Tr in Table 2) and accuracy is evaluated on another modality (denoted by Ev in Table 2). In retrieval tasks, we measure the accuracy of correct retrieval. For One-to-One case, we retrieve data sample in different modality by choosing the closest embedding from a single input embedding, while for Two-to-One case we choose the closest embedding from the centroid of two input embeddings in two modalities. We denote input and target modalities with $\rightarrow$ in Table 1.
@@
Table 1 shows the performance for one-to-one and two-to-one retrieval tasks. CentroBind consistently excels in one-to-one retrieval for text and audio modalities, while FABind performs better for video retrieval. This might be due to a strong dependency between text and video, which may be suitable for FABind anchored at text modality. A notable observation is that the centroid of video and audio embeddings achieves the best text retrieval performance including one-to-one retrieval. This implies complementary information exists and is captured by CentroBind.
Table 2: Accuracy results for Sarcasms and Speakers. (Tr: training modality, Ev: evaluation modality, $\mathcal{V}$: video, $\mathcal{A}$: audio, $\mathcal{T}$: text). Methods with asterisks denote accuracy evaluated in different experiment settings.
@@
Table 2 presents results for sarcasm detection and speaker classification tasks, where Sar-1 indicates Top-1 accuracy for sarcasm, and Spk-1, Spk-3, and Spk-5 represent Top-1, Top-3, and Top-5 accuracies for speaker classification. In this experiment, CentroBind consistently outperforms FABind across all pairs of train and evaluation modalities, which can be distributed to CentroBind generally learning a better shared embedding space than FABind.
Although a direct comparison is not feasible, we present the sarcasm detection accuracy of FactorCL (Liang et al., 2024b) and SimMMDG (Dong et al., 2023) approaches for reference. It is important to highlight that CentroBind and FABind are trained on a single modality (Tr) and evaluated on a different modality (Ev) in a zero-shot setting. In contrast, FactorCL and SimMMDG employ supervised contrastive learning, utilizing sarcasm labels during pretraining to capture task-specific features. At inference, all three modalities are simultaneously used to predict the class.
As analyzed in Section 2.3 and Section 3.2, these results highlight the CentroBind’s ability to preserve intra and shared information among modalities, which are useful in unknown downstream tasks. Moreover, the zero-shot setting verifies the multimodal alignment of both CentroBind and FABind.
@@ 5 Conclusions
In this paper, we mathematically analyze the limitations of fixed-anchor-bind methods (FABind), including over-reliance on the choice of anchor modality, and failing to capture both intra-information and shared information among non-anchored modalities. To overcome the over-reliance on an anchor modality, we propose CentroBind, which aligns multimodal embeddings to dynamic anchors constructed by centroids of the embeddings, hence removing the need for anchor modality. Moreover, we theoretically study CentroBind, showing that it captures intra- and shared information. Extensive experiments on both synthetic and real-world datasets show that CentroBind significantly outperforms FABind, providing a robust unified representation space and validating our analysis on CentroBind and FABind.
@@ Appendix A Related work
A.1 Multimodal alignment
@@
Multimodal learning addresses four key challenges (Liang et al., 2024c; Baltrušaitis et al., 2018; Liang et al., 2024d): managing interactions among redundant, unique, and synergistic features (Dumas et al., 2017; Liang et al., 2024a; b), aligning fine-grained and coarse-grained information (Wang et al., 2023; 2024a), reasoning across diverse features (Yang et al., 2023), and integrating external knowledge (Shen et al., 2022; Lyu et al., 2024). Among these challenges, multimodal alignment is one of the core challenges that many researchers aim to solve.
@@
A common method in multimodal alignment is using cross-modal alignment by using attention mechanisms between pairwise modalities, such as vision-language (Tan & Bansal, 2019) and vision-language-audio (Tsai et al., 2019). Another effective approach is leveraging graph neural networks to align multimodal datasets (Yang et al., 2021; Wilf et al., 2023). For instance, Yang et al. (2021) transforms unaligned multimodal sequence data into nodes, with edges capturing interactions across modalities over time. Wilf et al. (2023) build graph structures for each modality—visual, textual, and acoustic—and create edges to represent their interactions.
@@
To enhance the generalizability of cross-modal representations, Xia et al. (2024) employ a unified codebook approach, facilitating a joint embedding space for visual and audio modalities. Another prominent method (Radford et al., 2021) achieves cross-modal alignment by leveraging large collections of image-text pairs, making it a widely adopted strategy in multimodal learning (Zhang et al., 2022; Guzhov et al., 2022; Zhou et al., 2023).
A.2 Binding methods
@@
Recent studies have focused on aligning multimodal datasets by leveraging binding properties in various modalities. ImageBind (Girdhar et al., 2023) aligns multimodal data by using image representation as the anchor and aligning each modality’s embedding with the image embedding. Similarly, LanguageBind (Zhu et al., 2024) uses language representation as the anchor, aligning other modalities into the language space. PointBind (Guo et al., 2023) learns a joint embedding space across 3d point, language, image, and audio modalities by designating the point space as the central representation. Thanks to the efficacy of such a binding idea with a fixed anchor, several “-Bind” approaches have been studied in numerous domains (Teng et al., 2024; Xiao et al., 2024; Gao et al., 2024; Yang et al., 2024b; Balemans et al., 2024; Dhakal et al., 2024; Yang et al., 2024a) While these methods demonstrate strong performance in zero-shot cross-modality retrieval and classification tasks, they are constrained by their reliance on an existing single anchor modality.
@@
Several approaches have integrated additional knowledge into multimodal representation spaces to address this limitation. Freebind (Wang et al., 2024a) introduces bi-modality spaces to enhance a pretrained image-paired unified space. It generates pseudo-embedding pairs across diverse modality pairs and aligns them with the pre-trained unified space using contrastive learning. Omnibind (Wang et al., 2024b) leverages multiple pretrained multimodal models to construct pseudo item-pair retrievals based on top-1 recall across various modality combinations using pairwise cross-modal alignment. Both methods show promising results in cross-modal retrieval by incorporating extra spaces into existing pairwise binding spaces. However, they still rely on fixed (pre-trained) representation spaces.
@@
Unibind (Lyu et al., 2024) highlights the imbalanced representation when using image-centered representation spaces. To address this, Unibind employs large language models (LLMs) to create a unified and balanced representation space. It constructs a knowledge base with multimodal category descriptions, establishes LLM-augmented class-wise embedding centers, and aligns other modalities to these centers through contrastive learning. This approach attempts to balance representations across modalities but still depends heavily on large-scale pretrained LLMs and centers alignment around a single unified space.
@@ B.1 Proof of Proposition 1
Moreover, $f_{1}^{\rm suf}$ obtained in Definition 1 with proper choice of $\mathcal{Z}$ achieves the maximum mutual information, implying together with $I({\bf X};{\bf Y})\leq\min\{H({\bf X}),H({\bf Y})\}$ that $I(f_{1}^{\rm suf}({\bf X}_{1});{\bf X}_{1})=H({\bf X}_{1})$, where $H({\bf X}_{1})$ is the entropy of ${\bf X}_{1}$ (Polyanskiy & Wu, 2024). In other words, we have $H({\bf X}_{1}|f_{1}^{\rm suf}({\bf X}_{1}))=H({\bf X}_{1})-I(f_{1}^{\rm suf}({\bf X}_{1});{\bf X}_{1})=0$. This gives
@@
We conclude the proof of Proposition 1 by noting that the optimality of FABind (i.e., $I(f_{1}^{\rm suf}({\bf X}_{1});{\bf X}_{i})=I(f_{1}^{\rm suf}({\bf X}_{1});f_{i}^{\rm FB}({\bf X}_{i})),~{}\forall i\in\{2,\cdots,M\}$) yields
@@ B.2 Proof of Proposition 2
where the labeled inequalities follow from: $\rm(a)$ the non-negativity of mutual information; $\rm(b)$ the data processing inequality. This concludes the proof of Proposition 2.
B.3 Proof of Theorem 1
To prove Theorem 1, we leverage the reverse inequality of $M$-variable Hölder inequality (Seo, 2013, eq. (2.8)). For the sake of completeness, we state the inequality in Lemma 1.
Lemma 1 (Reverse inequality of the $M$-variable Hölder inequality (Seo, 2013)).
@@ B.3 Proof of Theorem 1
where the labeled (in)equalities follow from: $\rm(a)$ Lemma 1 and $C_{\mathcal{F},k,i}=\frac{(c_{\mathcal{F},k,i}^{\min}+c_{\mathcal{F},k,i}^{\max})^{2}}{4c_{\mathcal{F},k,i}^{\min}c_{\mathcal{F},k,i}^{\max}}$ with
@@
which concludes the proof of Theorem 1.
Appendix C Experiment details

### v2

@@ 4 Experiment
4.1 Experiments with synthetic datasets
Synthetic datasets.
We utilize a Gaussian mixture model (GMM) (Bishop & Nasrabadi, 2006) to generate synthetic multi-modal datasets. Let ${\bf X}_{i}$ denote the $i$-th modality dataset, with modal qualities arranged in ascending order based on their inherent correlation with the latent variable. Specifically, among $M$ modalities, ${\bf X}_{1}$ represents the least correlated modality, while ${\bf X}_{M}$ represents the most correlated. The class label, ${\bf Y}$, is assigned according to the GMM component identifier. A detailed description of the dataset generation process is provided in Appendix C.1.
@@ 4.1 Experiments with synthetic datasets
Figure 3 shows the classification accuracies with a synthetic dataset of $M=4$ modalities. To obtain the results in Figure 3(a), we initialize pretrained backbones for each modality, apply FABind (${\bf X}_{i}$-B) with anchor ${\bf X}_{i}$ or CentroBind (CB), and evaluate accuracy (acc(${\bf Z}_{i}$)) with embeddings from $i$-th modality. We provide acc(${\bf Z}_{i}$), without any binding, for a reference. Figure 3(a) verifies our analysis of FABind (Section 2.3) and CentroBind (Section 3.2): (1) the comparison between ${\bf X}_{1}$-B and ${\bf X}_{4}$-B shows the importance of choosing an anchor modality; (2) the comparison between acc(${\bf Z}_{4}$) and ${\bf X}_{1}$-B: acc(${\bf Z}_{4}$) shows a performance deterioration by FABind , demonstrating the impact on performance of FABind failure to capture intra information; (3) the proposed CB consistently outperforms FABind, indicating that CB successfully captures elements that FABind overlooks, including intra information and shared information among non-anchored modalities.
Figure 3(b) includes accuracies of FABind and CB with random backbones. Similarly, CB outperforms all baselines. Somewhat surprisingly, CB with random backbones (green curves) also performs better than FABind with pretrained backbones (red curves). This further supports our analysis that CentroBind is robust to backbone quality as it optimizes intra and shared information, whereas FABind is sensitive to the backbone quality. Overall, these empirical results validate our analysis of FABind and CentroBind. For clarity, we summarize the final accuracies in Table 3. We provide experimental results with $M=6,8$ in Appendix C.1. With the larger number of modalities, CB still outperforms the baselines.
Embedding visualization.
In addition, we visualize the embeddings learned by FABind and CentroBind using UMAP (McInnes et al., 2018) in Figure 4 (for details and visualizations using t-SNE (Van der Maaten & Hinton, 2008), see Appendix C.1). Figure 4 shows that CentroBind embeddings are better clustered, whereas FABind embeddings appear more scattered, implying that CentroBind yields better embeddings compared to FABind.
Stability.
We empirically examine the stability and convergence speed of both CentroBind and FABind. In Figure 7, we plot the training loss curves, which show similar behavior, suggesting that the adaptive anchor does not introduce issues related to convergence or stability. Detailed discussions can be found in Appendix C.1.
(a) FABind.
(b) CentroBind.
Figure 4: Representation visualization via UMAP.
@@ 4.2 Experiments with real-world dataset
In this section, we provide experiment results with a real-world dataset. We compare CentroBind, FABind anchored at text modality, UniBind (Lyu et al., 2024), AudioCLIP (Guzhov et al., 2022), and ViT-Lens (Lei et al., 2024) (see implementation detail in Appendix C.2). We utilize the MUStARD (Castro et al., 2019) and DreamBooth dataset (Ruiz et al., 2023). [^4] For experiments on the DreamBooth dataset, please refer to Appendix C.3. The MUStARD dataset consists of 690 video, audio and text for sarcasm detection with labels such as sarcasm indicators and speaker names. For the backbones in FABind and CentroBind , we use the pretrained VideoMAE (Tong et al., 2022) for video data, the pretrained WaveLM (Chen et al., 2022) for audio data, and the pretrained BERT (Devlin et al., 2019) for text data. A detailed description of the training setting is provided in Appendix C.2.
@@
We perform evaluations in zero-shot binary and multi-class classification tasks, One-to-One cross-modal retrieval, and Two-to-One cross-modal retrieval. For classification tasks, we use a Multi-Layer Perceptron (MLP) to perform sarcasm detection as a binary classification and speaker classification with $23$ multi-class categories. In particular, MLP is trained on embeddings in a single modality (denoted by Tr in Table 2) and accuracy is evaluated on another modality (denoted by Ev in Table 2). In retrieval tasks, we measure the accuracy of correct retrieval. For the One-to-One case, we retrieve data sample in different modality by choosing the closest embedding from a single input embedding, while for the Two-to-One case we choose the closest embedding from the centroid of two input embeddings in two modalities. We denote input and target modalities with $\rightarrow$ in Table 1.
@@
Table 1 shows the performance for One-to-One and Two-to-One retrieval tasks. CentroBind consistently excels in One-to-One retrieval for text and audio modalities, while FABind performs better for video retrieval. This might be due to the power of text to describe, which may be suitable for FABind anchored at text modality. A notable observation is that the centroid of video and audio embeddings achieves the best text retrieval performance. This implies complementary information exists and is captured by CentroBind.
Table 2: Accuracy results for Sarcasms and Speakers. ($\mathcal{V}$: video, $\mathcal{A}$: audio, $\mathcal{T}$: text). Asterisks denote accuracy evaluated in different settings.
@@
Table 2 presents results for sarcasm detection and speaker classification tasks, where Sar-1 indicates Top-1 accuracy for sarcasm, and Spk-$k$, $k=1,3,5$ represent Top-k accuracies for speaker classification. It is important to highlight that CentroBind and FABind are trained on a single modality (Tr) and evaluated on a different modality (Ev) in a zero-shot setting, which can effectively measure the ability of multi-modal alignment. In this experiment, CentroBind consistently outperforms FABind and UniBind across all pairs of train and evaluation modalities, which can be distributed to CentroBind generally learning a better unified embedding space than FABind. UniBind performs poorly in the zero-shot cross-modal experiment, which we believe is due to its insufficient multi-modal alignment. Since UniBind utilizes LLM-augmented descriptions for each modality and binds other encoders to these descriptions, multi-modal alignment may fail if the descriptions are dispersed across the embedding space. As analyzed in Section 2.3 and Section 3.2, these results highlight the CentroBind’s ability to preserve intra and shared information among modalities, which are useful in unknown downstream tasks. Moreover, the zero-shot setting verifies the multi-modal alignment of CentroBind.
@@ 5 Conclusions
In this paper, we analyze the limitations of fixed-anchor-binding (FABind) methods, including their over-reliance on a single anchor modality and their inability to capture both intra-modal and shared information among non-anchored modalities. To address these shortcomings, we propose adaptive anchor binding methods such as CentroBind, which align multi-modal embeddings to centroid-based anchors, removing the need for a fixed anchor modality. Our approach generalizes and extends methods like ImageBind. We also provide a theoretical analysis showing that CentroBind effectively captures both intra-modal and shared inter-modal information. Experiments on synthetic and real-world datasets demonstrate that CentroBind significantly outperforms FABind, offering a robust unified representation space and validating our theoretical insights.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
@@ Appendix A Related work
A.1 Multi-modal learning
Multi-modal learning has gained significant attention in recent years due to its potential to enhance machine learning models by leveraging diverse data modalities, such as text, images, audio, and video. By combining these modalities, multi-modal learning seeks to mimic human-like perception, thereby improving performance across a wide range of applications, from healthcare to natural language processing. Common supervised multi-modal learning tasks include audio-visual classification (Peng et al., 2022; Feichtenhofer et al., 2019; Zhu & Rahtu, 2022), visual question answering (Antol et al., 2015; Guo et al., 2021), and vision-language tasks (Xu et al., 2015; Radford et al., 2021), as well as vision-audio-language tasks (Aytar et al., 2017; Harwath et al., 2018).
Typically, these models integrate uni-modal features extracted by modality-specific encoders (Seichter et al., 2021; Nagrani et al., 2021; Wu et al., 2022; Wang et al., 2020; Peng et al., 2022). For instance, Madaan et al. (2024) introduce inter- and intra-modality modeling frameworks that treat the target as a composition of multiple modalities. Similarly, Du et al. (2023) propose a late-fusion approach for supervised multi-modal tasks, demonstrating that insufficient feature extraction from individual modalities negatively affects the model’s generalization ability. Additionally, Zhang et al. (2024) address joint optimization by alternating between uni-modal learning scenarios and integrating modality-specific encoders with a unified head shared across all modalities.
Multi-view learning is closely related to multi-modal learning. Early work by Tian et al. (2020a) study multi-view representation learning through predictive and contrastive approaches across multiple modalities, demonstrating the efficacy of contrastive learning in multi-view settings. Building on these insights, numerous subsequent studies have integrated contrastive learning into multi-view tasks. For example, Xing et al. (2019) propose an approach that adaptively combines image-text representations for few-shot learning; however, this method still relies on labeled data. Some research leverages anchors or a fused modality as the anchor representation. Zeng et al. (2021) introduce a unified prototype representation to address cross-modal retrieval imbalance by employing a reconstruction loss, but its reliance on unified prototypes as anchors restricts the capture of inter-view information across modalities. Meanwhile, Huang et al. (2023) develop a tri-modal alignment pre-training task—extending text-video alignment to include a fused modality using pairwise contrastive learning; however, it does not explicitly handle intra-view learning. Furthermore, Wang et al. (2022b) present a cross-modal data augmentation technique for image-text multi-view learning, randomly replacing visually grounded words with diverse image patches to increase data variety and encourage token-level interaction across modalities. Nevertheless, this approach still overlooks intra-view information, indicating the need for methods that jointly consider both inter- and intra-view representations.
A.2 Multi-modal alignment
@@
Multi-modal learning addresses four key challenges (Liang et al., 2024c; Baltrušaitis et al., 2018; Liang et al., 2024d): managing interactions among redundant, unique, and synergistic features (Dumas et al., 2017; Liang et al., 2024a, b), aligning fine-grained and coarse-grained information (Wang et al., 2023, 2024a), reasoning across diverse features (Yang et al., 2023), and integrating external knowledge (Shen et al., 2022; Lyu et al., 2024). Among these challenges, multi-modal alignment is one of the core challenges that many researchers aim to solve.
@@
A common method in multi-modal alignment is using cross-modal alignment by using attention mechanisms between pairwise modalities, such as vision-language (Tan & Bansal, 2019) and vision-language-audio (Tsai et al., 2019). Another effective approach is leveraging graph neural networks to align multi-modal datasets (Yang et al., 2021; Wilf et al., 2023). For instance, Yang et al. (2021) transform unaligned multi-modal sequence data into nodes, with edges capturing interactions across modalities over time. Wilf et al. (2023) build graph structures for each modality—visual, textual, and acoustic—and create edges to represent their interactions.
@@
To enhance the generalizability of cross-modal representations, Xia et al. (2024) employ a unified codebook approach, facilitating a joint embedding space for visual and audio modalities. Another prominent method achieves cross-modal alignment by leveraging large collections of image-text pairs, making it a widely adopted strategy in multi-modal learning (Radford et al., 2021; Zhang et al., 2022; Guzhov et al., 2022; Zhou et al., 2023).
A.3 Binding methods
@@
Recent studies have focused on aligning multi-modal datasets by leveraging binding properties in various modalities. ImageBind (Girdhar et al., 2023) align multi-modal data by using image representation as the anchor and aligning each modality embedding with the image embedding. Similarly, LanguageBind (Zhu et al., 2024) use language representation as the anchor, aligning other modalities into the language space. PointBind (Guo et al., 2023) learn a joint embedding space across 3d point, language, image, and audio modalities by designating the point space as the central representation. Thanks to the efficacy of such a binding idea with a fixed anchor, several “-Bind” approaches have been studied in numerous domains (Teng et al., 2024; Xiao et al., 2024; Gao et al., 2024; Yang et al., 2024b; Balemans et al., 2024; Dhakal et al., 2024; Yang et al., 2024a) While these methods demonstrate strong performance in zero-shot cross-modality retrieval and classification tasks, they are constrained by their reliance on an existing single anchor modality.
@@
Several approaches have integrated additional knowledge into multi-modal representation spaces to address this limitation. Freebind (Wang et al., 2024a) introduce bi-modality spaces to enhance a pretrained image-paired unified space. It generates pseudo-embedding pairs across diverse modality pairs and aligns them with the pre-trained unified space using contrastive learning. Omnibind (Wang et al., 2024b) leverage multiple pretrained multi-modal models to construct pseudo item-pair retrievals based on top-1 recall across various modality combinations using pairwise cross-modal alignment. Both methods show promising results in cross-modal retrieval by incorporating extra spaces into existing pairwise binding spaces. However, they still rely on fixed (pre-trained) representation spaces.
@@
Unibind (Lyu et al., 2024) highlight the imbalanced representation when using image-centered representation spaces. To address this, Unibind employs large language models (LLMs) to create a unified and balanced representation space. It constructs a knowledge base with multi-modal category descriptions, establishes LLM-augmented class-wise embedding centers, and aligns other modalities to these centers through contrastive learning. This approach attempts to balance representations across modalities but still depends heavily on large-scale pretrained LLMs and centers alignment around a single unified space, namely, text (language).
ViT-Lens (Lei et al., 2024) build upon the Vision Transformer (ViT) (Dosovitskiy et al., 2021) and multi-modal foundational models like CLIP (Radford et al., 2021) to align multiple modalities. It extends ViT by incorporating an additional embedding layer and attention layer for each modality, which are trained via contrastive learning involving embeddings generated by the CLIP and the ViT models. This approach generalizes FABind by allowing more than one fixed anchor modality; specifically, image and text in this case. CentroBind could also adopt a similar strategy, leveraging the powerful ViT model for modality alignment while adaptively computing anchors based on their centroids.
@@ Appendix B Proofs
Moreover, $f_{1}^{\rm suf}$ obtained in Definition 2.1 with proper choice of $\mathcal{Z}$ achieves the maximum mutual information, implying together with $I({\bf X};{\bf Y})\leq\min\{H({\bf X}),H({\bf Y})\}$ that $I(f_{1}^{\rm suf}({\bf X}_{1});{\bf X}_{1})=H({\bf X}_{1})$, where $H({\bf X}_{1})$ is the entropy of ${\bf X}_{1}$ (Polyanskiy & Wu, 2024). In other words, we have $H({\bf X}_{1}|f_{1}^{\rm suf}({\bf X}_{1}))=H({\bf X}_{1})-I(f_{1}^{\rm suf}({\bf X}_{1});{\bf X}_{1})=0$. This gives
@@
We conclude the proof of Proposition 2.2 by noting that the optimality of FABind (i.e., $I(f_{1}^{\rm suf}({\bf X}_{1});{\bf X}_{i})=I(f_{1}^{\rm suf}({\bf X}_{1});f_{i}^{\rm FB}({\bf X}_{i})),~{}\forall i\in\{2,\cdots,M\}$) yields
@@
where the labeled inequalities follow from: $\rm(a)$ the non-negativity of mutual information; $\rm(b)$ the data processing inequality. This concludes the proof of Proposition 2.3.
B.3 Proof of Theorem 3.1
To prove Theorem 3.1, we leverage the reverse inequality of $M$-variable Hölder inequality (Seo, 2013, eq. (2.8)). For the sake of completeness, we state the inequality in Lemma B.1.
Lemma B.1 (Reverse inequality of the $M$-variable Hölder inequality (Seo, 2013)).
@@
where the labeled (in)equalities follow from: $\rm(a)$ Lemma B.1 and $C_{\mathcal{F},k,i}=\frac{(c_{\mathcal{F},k,i}^{\min}+c_{\mathcal{F},k,i}^{\max})^{2}}{4c_{\mathcal{F},k,i}^{\min}c_{\mathcal{F},k,i}^{\max}}$ with
@@
which concludes the proof of Theorem 3.1.
Appendix C Experiment